If you need to buy 1,000,000 in your local currency for business when your bank provides...

  1. If you need to buy 1,000,000 in your local currency for business when your bank provides the following bid-ask rates: 1.3895/USD-1.3905/USD, what will be the dollar value after purchase?

CAN YOU SHOW STEPS
Answer #1

Buy rate per $ rate = 1.3905

We need to convert this into per local Rate

Local Rate = 1/dollar rate = 1 / 1.3905 = 0.7192 / $

Dollar value =Amount in local currency * Local rate

Dollar value = 10,00,000 * 0.7192 = $719200

Dollar value after purchase = $719200

Note: Sale rate = 1.3895 / $ & Buy rate = 1.3905 / $

As we have to buy local currency so we will choose buy rate for local currency.  Here rate given is $ rate which need to convert into local currency rate by dividing with 1.